# Birational Geometry of Special Quotient Foliations and Chazy’s Equations

### v1

@@ 1. Introduction and results
it is tangent to the nodal cubic $03xy^{2}-y^{3}-3xyz+xz^{2}=$ 0, which gives rise to the link, and to its inflectional lines $x=0$ and $x-3y+z=0$.
@@
It is tangent to the nodal cubic $02y^{3}-2xy^{2}+2xyz-xz^{2}=$ 0, which gives rise to the link, to its inflectional line $x=0$, and to the smooth conic $y^{2}+4xy-2xz-x^{2}=0$. It is the only degree-two foliation on $\mathbb{P}^{2}$ simultaneously tangent to these cubic, conic and line.
@@
It is tangent to the nodal cubic $03y^{3}-xy^{2}+xyz-xz^{2}=$ 0, which gives rise to the link, to its inflectional line $x=0$, and to the nodal cubic $08y^{3}-15xy^{2}+6xyz-3xz^{2}-6x^{2}y+6x^{2}z+x^{3}=$ 0. It is the only degree-two foliation on $\mathbb{P}^{2}$ tangent to both cubics.
@@
(see [Cha11, p. 336]; for their integration, see [Cha11, p. 343], [Cos00, Sections 6.4–6.6], or [Gui12]). These equations have the form $x^{\prime\prime\prime}=P(x,x^{\prime},x^{\prime\prime})$, with $P(x,y,z)$ a polynomial which is of degree $4$ when $x$, $y$ and $z$ are, respectively, given the weights $1$, $2$ and $3$. They may be described by polynomial vector fields on $\mathbb{C}^{3}$ of the form
@@
The previous quasihomogeneity property for $P$ is equivalent to the fact that $P(\lambda x,\lambda^{2}y,\lambda^{3}z)=\lambda^{4}P(x,y,z)$. The transformation (7) acts upon a vector field $W$ as above by dividing it by $\lambda$. The action preserves thus the foliation on $\mathbb{C}^{3}$ induced by $W$, and induces a foliation on the quotient of $\mathbb{C}^{3}\setminus\{0\}$ under the above action, the two-dimensional variety known as the weighted projective plane $\mathbb{P}(1,2,3)$ (to be discussed in Section 2.3).
@@
Brunella considered that “it would be natural to look for other types of birational models [for $\mathcal{F}_{3}$]” [Bru15, p. 54]. The above result provides alternative models, not just for $\mathcal{F}_{3}$, but for $\mathcal{F}_{4}$ and $\mathcal{F}_{6}$ as well. It will follow, on the one hand, from the definitions of these foliations as quotients, and, on the other, from the description of the foliations induced by the corresponding Chazy equations appearing in [Gui12, pp. 71–74]. We will revisit this last result in Section 4. Theorems 2 and 3, and an alternative proof of Theorem 1, will follow from it, and from an explicit birational equivalence between $\mathbb{P}(1,2,3)$ and $\mathbb{P}^{2}$.
@@
For the models for $\mathcal{F}_{3}$ and $\mathcal{F}_{4}$ given in Theorems 1 and 2, Propositions 13 and 15 will give explicit formulas for generators of these groups.
@@
Its set of invariant algebraic curves is composed by the nodal cubic $xy^{2}+yz^{2}+zx^{2}-3xyz=0$, representing the link, and by the three coordinate lines, tangent to the latter. It is the only degree-two foliation of $\mathbb{P}^{2}$ leaving invariant this configuration of curves. Its group of birational automorphisms is generated by the quartic de Jonquières involution
@@
Let $\lambda\in\mathbb{C}\setminus\{0,1\}$. The quotient of the degree-one foliation on $\mathbb{P}^{2}$ given by $\lambda YZ\,\mathrm{d}X-XZ\,\mathrm{d}Y+(1-\lambda)XY\,\mathrm{d}Z=0$, under the action of the standard quadratic Cremona transformation $(X:Y:Z)\mapsto(YZ:ZX:XY)$, is birationally equivalent to the foliation $\mathcal{G}_{\lambda}$ of degree three on $\mathbb{P}^{2}$ given by
@@
The article is organized in the following way. After reviewing some background material in Section 2, we recall the definition of the three special quotient foliations in Section 3. In Section 4 we study the relations between the Chazy equations and the special quotient foliations, and establish Theorem 4. This will allow us to give an alternative proof of Theorem 1 in Section 4.1.1, and to prove Theorems 2 and 3 in Sections 4.2.1 and 4.3.1, respectively. Theorems 6, 7 and 8 will be established in the same section. Section 5 will be devoted to Theorem 9, and Theorem 10 will proved in Section 6. Theorem 5, whose proof is more analytic in nature, will be established in Section 7, where we will also calculate the groups of birational automorphisms of hyperbolic linear foliations of the plane (Theorem 17).
Acknowledgments
@@ 2.1. On singular holomorphic foliations
where $a_{i},b_{i}\in\mathcal{O}(U_{i})$, with $\gcd(a_{i},b_{i})=1$, such that, along $U_{i}\cap U_{j}\neq\emptyset$, $\omega_{i}=g_{ij}\,\omega_{i}$ for $g_{ij}\in\mathcal{O}^{*}(U_{i}\cap U_{j})$. The singularities of the foliation are given by the zeros of the forms $\omega_{i}$. The conditions $\gcd(a_{i},b_{i})=1$ ensure that the singular set is locally finite. The foliation locally defined by the $1$-form $\omega=a(x,y)\,\mathrm{d}x+b(x,y)\,\mathrm{d}y$ may also be defined by its dual vector field $v=b(x,y)\,\partial/\partial x-a(x,y)\,\partial/\partial y$. A singular point of the foliation, a point where both $a$ and $b$ vanish, is said to be non-degenerate if the eigenvalues of the linear part of $v$ at this singular point are both non-zero. In such a case, if $\lambda_{1}$ and $\lambda_{2}$ are these eigenvalues, the eigenvalues of the singular point are said to be $\lambda_{2}:\lambda_{1}$ (or $\lambda_{1}:\lambda_{2}$). A singularity $p$ of the foliation is said to be reduced if $v$ has a non-nilpotent linear part at $p$, and if it has either one vanishing eigenvalue, or it is non-degenerate and $\lambda_{2}/\lambda_{1}\not\in\mathbb{Q}^{+}$. After a finite number of blow-ups, every singularity of a foliation is replaced by finitely many reduced singularities along the exceptional divisor (Seidenberg’s reduction of singularities).
The non-degenerate singularities of eigenvalues $n:1$, with $n$ a strictly positive integer, admit the local Poincaré-Dulac normal form, $(ny+\epsilon x^{n})\,\mathrm{d}x-x\,\mathrm{d}y$, with $\epsilon\in\{0,1\}$, for which the curve $x=0$ is invariant. For $\epsilon=1$, this is the only invariant curve. In particular, if there is more than one invariant curve through such a singularity, it is linearizable (has $\epsilon=0$ in its normal form). For $\epsilon=0$, we have the first integral $y/x^{n}$, and, with it, the invariant curves of the form $y=cx^{n}$, any two of which have a contact of order $n$. When $n>1$, blowing up the foliation $ny\,\mathrm{d}x-x\,\mathrm{d}y$ produces an invariant divisor with two singularities, one of type $n:1-n$, and one of type $n-1:1$, a linearizable one. Blowing up the radial foliation $y\,\mathrm{d}x-x\,\mathrm{d}y$ (the above foliation in the case $n=1$) produces a dicritical exceptional divisor, one that is everywhere transverse to the foliation.
@@
Let $Q(x:y:z)=(yz:xz:xy)$ be the standard quadratic Cremona map. Let $p_{1}$, $p_{2}$ and $p_{3}$ be the vertices of the coordinate triangle $xyz=0$. Let $\mathcal{F}$ be a foliation of the plane of degree $\deg(\mathcal{F})$, with $l(p_{i},\mathcal{F})\geq 0$. Let $\overline{\mathcal{F}}$ be the transformed foliation of $\mathcal{F}$ under $Q$ (with finite singular set). Then,
@@ 2.2. The Klein surface singularities of type $A_{n}$
We follow [dlHS79, Section IV] for the discussion that follows. Let $n\geq 1$. Consider the analytic space
@@
Let $C_{n+1}\subset\mathrm{GL}(2,\mathbb{C})$ be the subgroup generated by $(s,t)\mapsto(\beta^{-1}s,\beta t)$, with $\beta$ a primitive $(n+1)$-th root of unity. Let $X_{n+1}=\mathbb{C}^{2}/C_{n+1}$. The injective analytic map $\phi_{n+1}:X_{n+1}\to A_{n}$ given by $\phi_{n+1}(s,t)=(s^{n+1},t^{n+1},st)$, realizes an analytic equivalence between $X_{n+1}/C_{n+1}$ and $A_{n}$.
For the desingularization of $A_{n}$, consider $n+1$ copies $R_{0},\ldots,R_{n}$ of $\mathbb{C}^{2}$, with coordinates $(u_{i},v_{i})$ on $R_{i}$, glued by the functions $\varphi_{k-1}:R_{k-1}\dashrightarrow R_{k}$ given by
@@
define a global map $\rho:M_{n}\to A_{n}$, the resolution of $A_{n}$.
@@
Holomorphic actions of finite groups are holomorphically linearizable in a neighborhood a fixed point, and thus $A_{n}$ gives the local model for the quotient of the action of $\mathbb{Z}_{n+1}$ generated by a transformation that, at a fixed point, has eigenvalues $\beta$ and $\beta^{n}$ [BHPV04, Thm. 5.4, Ch. III].
@@ 2.3. A weighted projective plane and the standard one
The quotient of $\mathbb{C}^{3}\setminus\{0\}$ under the action of the weighted homotheties (7) is the weighted projective plane $\mathbb{P}(1,2,3)$. The class of $(x,y,z)\in\mathbb{C}^{3}\setminus\{0\}$ in $\mathbb{P}(1,2,3)$ will be denoted by $[x:y:z]$.111 We call attention for the use, throughout the article, of the notation $[x:y:z]$ for points in the weighted projective plane $\mathbb{P}(1,2,3)$, and of $(x:y:z)$ for points in the standard projective plane $\mathbb{P}^{2}$. The weighted projective plane $\mathbb{P}(1,2,3)$ is covered by three charts, each one of which is either $\mathbb{C}^{2}$, or its quotient under the action of a finite linear group:
@@
$j([x:y:z])=j\left(\left[1:\frac{y}{x^{2}},\frac{z}{x^{3}}\right]\right)=\left(1:\frac{y}{x^{2}},\frac{z}{x^{3}}\right)=(x^{3}:xy:x^{2}z),$
@@
(14) $[x:y:z]\mapsto(x^{3}:xy:z),$
having inverse $(X:Y:Z)\mapsto[X:XY:X^{2}Z]$. A factorization of this birational map is given as follows. Consider the resolution $\pi:S\to\mathbb{P}(1,2,3)$ of $\mathbb{P}(1,2,3)$, obtained by desingularizing the $A_{1}$ and $A_{2}$ singular points, as explained in Section 2.2. The surface $S$ has a $(-2)$-curve $C_{1}$ (i.e. a smooth rational curve of self-intersection $-2$), corresponding to the resolution of $A_{1}$, and a chain of two $(-2)$-curves, $D_{1}$ and $D_{2}$, intersecting transversely at one point, corresponding to the resolution of $A_{2}$. The strict transform of the curve $\mathcal{\ell}$ given by $x=0$ is a rational curve $\overline{\ell}$ in $S$ of self-intersection $-1$, intersecting $C_{1}$ and $D_{1}$ at one point each, transversely. The contractions of $\overline{\ell}$ and the transforms of $D_{1}$ and $D_{2}$, in this order, produce the projective plane, where the transform of $C_{1}$ is a straight line. See Figure 2.
@@ 3. Three special quotient foliations
For $n\in\{3,4,6\}$, the foliation $\mathcal{F}_{n}$ admits a simple description as quotient of a foliation on a rational surface under the action of a cyclic group of automorphisms of order $n$. This surface has an invariant cycle of rational curves $\Delta_{n}$ of length $n$, whose components are cyclically permuted by the automorphism, and which produces the link in the quotient. This is schematically presented in Figure 3. We now present these foliations and exhibit some of their birational symmetries (Theorem 5 will establish that there are no further ones).

### v2

@@ Abstract.
Published as: Guillot, A. and Mendes, L. G. Birational Geometry of Special Quotient Foliations and Chazy’s Equations, Bull. Sci. Math. 209 (2026), art. no. 103792, doi:10.1016/j.bulsci.2025.103792
Γ Δ  $C$ BY 4.0. This work is licensed under a Creative Commons Attribution 4.0 License
@@ 1. Introduction and results
it is tangent to the nodal cubic $3xy^{2}-y^{3}-3xyz+xz^{2}=0$, which gives rise to the link, and to its inflectional lines $x=0$ and x3y+z=0x-3y+z=0.
@@
It is tangent to the nodal cubic $2y^{3}-2xy^{2}+2xyz-xz^{2}=0$, which gives rise to the link, to its inflectional line $x=0$, and to the smooth conic $y^{2}+4xy-2xz-x^{2}=0$. It is the only degree-two foliation on $\mathbb{P}^{2}$ simultaneously tangent to these cubic, conic and line.
@@
It is tangent to the nodal cubic $3y^{3}-xy^{2}+xyz-xz^{2}=0$, which gives rise to the link, to its inflectional line $x=0$, and to the nodal cubic $8y^{3}-15xy^{2}+6xyz-3xz^{2}-6x^{2}y+6x^{2}z+x^{3}=0$. It is the only degree-two foliation on $\mathbb{P}^{2}$ tangent to both cubics.
@@
(see [Cha11, p. 336]; for their integration, see [Cha11, p. 343], [Cos00, Sections 6.4–6.6], or [Gui12]). These equations have the form $x^{\prime\prime\prime}=P(x,x^{\prime},x^{\prime\prime})$, with $P(x,y,z)$ a polynomial which is of degree $4$ when $x$, $y$ and $z$ are, respectively, given the weights $1$, 22 and $3$. They may be described by polynomial vector fields on $\mathbb{C}^{3}$ of the form
@@
The previous quasihomogeneity property for $P$ is equivalent to the fact that $P(\lambda x,\lambda^{2}y,\lambda^{3}z)=\lambda^{4}P(x,y,z)$. The transformation (7) acts upon a vector field $W$ as above by dividing it by $\lambda$. The action preserves thus the foliation on $\mathbb{C}^{3}$ induced by $W$, and induces a foliation on the quotient of $\mathbb{C}^{3}\setminus\{0\}$ under the action (7), the two-dimensional variety known as the weighted projective plane $\mathbb{P}(1,2,3)$ (to be discussed in Section 2.3).
@@
Brunella considered that “it would be natural to look for other types of birational models [for $\mathcal{F}_{3}$]” [Bru15, p. 54]. The above result provides alternative models, not just for $\mathcal{F}_{3}$, but for $\mathcal{F}_{4}$ and $\mathcal{F}_{6}$ as well. It will follow, on the one hand, from the definitions of these foliations as quotients, and, on the other, from the description of the foliations induced by the corresponding Chazy equations appearing in [Gui12, pp. 71–74]. We will revisit this last result in Section 4. Theorems 2 and 3, and an alternative proof of Theorem 1, will follow from it, and from a particular explicit birational equivalence between $\mathbb{P}(1,2,3)$ and $\mathbb{P}^{2}$.
@@
For the models for $\mathcal{F}_{3}$ and $\mathcal{F}_{4}$ given in Theorems 1 and 2, Propositions 15 and 17 will give explicit formulas for generators of these groups.
@@
Its set of invariant algebraic curves is composed by the nodal cubic $xy^{2}+yz^{2}+zx^{2}-3xyz=0$, representing the link, and by the three coordinate lines, which are tangent to the latter. It is the only degree-two foliation of $\mathbb{P}^{2}$ leaving invariant this configuration of curves. Its group of birational automorphisms is generated by the quartic de Jonquières involution
@@
Let $\lambda\in\mathbb{C}\setminus\{0,1\}$. The quotient of the degree-one foliation on $\mathbb{P}^{2}$ given by $\lambda YZ\,\mathrm{d}X-XZ\,\mathrm{d}Y+(1-\lambda)XY\,\mathrm{d}Z=0$ under the action of the standard quadratic Cremona transformation $(X:Y:Z)\mapsto(YZ:ZX:XY)$, is birationally equivalent to the foliation $\mathcal{G}_{\lambda}$ of degree three on $\mathbb{P}^{2}$ given by
@@
The article is organized in the following way. After reviewing some background material in Section 2, we recall the definition of the three special quotient foliations in Section 3. In Section 4 we study the relations between the Chazy equations and the special quotient foliations, and establish Theorem 4. This will allow us to give an alternative proof of Theorem 1 in Section 4.1.4, and to prove Theorems 2 and 3 in Sections 4.2.4 and 4.3.4, respectively. Theorems 6, 7 and 8 will be established in the same section. Section 5 will be devoted to Theorem 9, and Theorem 10 will be proved in Section 6. Theorem 5, whose proof is more analytic in nature, will be established in Section 7, where we will also calculate the groups of birational automorphisms of hyperbolic linear foliations of the plane (Theorem 21).
@@ 2.1. On singular holomorphic foliations
where $a_{i},b_{i}\in\mathcal{O}(U_{i})$, with $\gcd(a_{i},b_{i})=1$, such that, along $U_{i}\cap U_{j}\neq\emptyset$, $\omega_{i}=g_{ij}\,\omega_{i}$ for $g_{ij}\in\mathcal{O}^{*}(U_{i}\cap U_{j})$. The singularities of the foliation are given by the zeros of the forms $\omega_{i}$. The conditions $\gcd(a_{i},b_{i})=1$ ensure that the singular set is locally finite. The foliation locally defined by the $1$-form $\omega=a(x,y)\,\mathrm{d}x+b(x,y)\,\mathrm{d}y$ may also be defined by its dual vector field $v=b(x,y)\,\partial/\partial x-a(x,y)\,\partial/\partial y$. A singular point of the foliation, a point where both $a$ and $b$ vanish, is said to be non-degenerate if the eigenvalues of the linear part of $v$ at this singular point are both non-zero. In such a case, if $\lambda_{1}$ and $\lambda_{2}$ are these eigenvalues, the eigenvalues of the singular point are said to be $\lambda_{2}:\lambda_{1}$ (or $\lambda_{1}:\lambda_{2}$). A singularity $p$ of the foliation is said to be reduced if $v$ has a non-nilpotent linear part at $p$, and if it either has one vanishing eigenvalue, or is non-degenerate and $\lambda_{2}/\lambda_{1}\not\in\mathbb{Q}^{+}$. After a finite number of blow-ups, every singularity of a foliation is replaced by finitely many reduced singularities along the exceptional divisor (Seidenberg’s reduction of singularities).
The non-degenerate singularities of eigenvalues $n:1$, with $n$ a strictly positive integer, admit the local Poincaré-Dulac normal form, $(ny+\epsilon x^{n})\,\mathrm{d}x-x\,\mathrm{d}y$, with $\epsilon\in\{0,1\}$, for which the curve $x=0$ is invariant. For $\epsilon=1$, this is the only invariant curve. In particular, if there is more than one invariant curve through such a singularity, it is linearizable (has $\epsilon=0$ in its normal form). For $\epsilon=0$, we have the first integral $y/x^{n}$, and, with it, the invariant curves of the form $y=cx^{n}$, any two of which have a contact of order $n$. When $n>1$, blowing up the foliation $ny\,\mathrm{d}x-x\,\mathrm{d}y$ produces an invariant divisor with two singularities, one of type n:1nn:1-n, and a linearizable one of type n1:1n-1:1. Blowing up the radial foliation $y\,\mathrm{d}x-x\,\mathrm{d}y$ (the above foliation in the case $n=1$) produces a dicritical exceptional divisor, one that is everywhere transverse to the foliation.
@@
Let $Q(x:y:z)=(yz:xz:xy)$ be the standard quadratic Cremona map. Let $p_{1}$, $p_{2}$ and $p_{3}$ be the vertices of the coordinate triangle $xyz=0$. Let $\mathcal{F}$ be a foliation of the plane of degree $\deg(\mathcal{F})$, with $l(p_{i},\mathcal{F})\geq 0$ for every $i$. Let $\overline{\mathcal{F}}$ be the transformed foliation of $\mathcal{F}$ under $Q$ (with finite singular set). Then,
@@ 2.2. The Klein surface singularities of type $A_{n}$
We follow [dlHS79, Sect. IV] for the discussion that follows. Let $n\geq 1$. Consider the analytic space
@@
Let $\beta$ be a primitive $(n+1)$-th root of unity, and let $C_{n+1}\subset\mathrm{GL}(2,\mathbb{C})$ be the group generated by $(s,t)\mapsto(\beta^{-1}s,\beta t)$. The analytic map $\phi_{n+1}:\mathbb{C}^{2}/C_{n+1}\to A_{n}$ given by $\phi_{n+1}(s,t)=(s^{n+1},t^{n+1},st)$ realizes an analytic equivalence between $\mathbb{C}^{2}/C_{n+1}$ and $A_{n}$.
The minimal desingularization of $A_{n}$ may be given as follows. Consider $n+1$ copies $R_{0},\ldots,R_{n}$ of $\mathbb{C}^{2}$, with coordinates $(u_{i},v_{i})$ on $R_{i}$, glued by the functions $\varphi_{k-1}:R_{k-1}\dashrightarrow R_{k}$ given by
@@
define a global map $\rho:M_{n}\to A_{n}$, the minimal resolution of $A_{n}$.
@@
Holomorphic actions of finite groups are holomorphically linearizable in a neighborhood of a fixed point, and thus $A_{n}$ gives the local model for the quotient of the action of $\mathbb{Z}_{n+1}$ generated by a transformation that, at a fixed point, has eigenvalues $\beta$ and $\beta^{n}$ [BHPV04, Thm. 5.4, Ch. III].
@@ 2.3. A weighted projective plane and the standard one
The quotient of $\mathbb{C}^{3}\setminus\{0\}$ under the action of the weighted homotheties (7) is the weighted projective plane $\mathbb{P}(1,2,3)$. The class of $(x,y,z)\in\mathbb{C}^{3}\setminus\{0\}$ in $\mathbb{P}(1,2,3)$ will be denoted by [x:y:z][x:y:z]. [^1] We draw the reader’s attention towards the systematic use, throughout this article, of the notation $[x:y:z]$ for points in the weighted projective plane $\mathbb{P}(1,2,3)$, and of $(x:y:z)$ for points in the standard projective plane $\mathbb{P}^{2}$. The weighted projective plane $\mathbb{P}(1,2,3)$ is covered by three charts, each one of which is either $\mathbb{C}^{2}$, or its quotient under the action of a finite linear group:
@@
$j([x:y:z])=j\left(\left[1:\frac{y}{x^{2}}:\frac{z}{x^{3}}\right]\right)=\left(1:\frac{y}{x^{2}}:\frac{z}{x^{3}}\right)=(x^{3}:xy:x^{2}z),$
@@
(14) $x:y:z]\mapsto(x^{3}:xy:z),$
having inverse $(X:Y:Z)\mapsto[X:XY:X^{2}Z]$. A factorization of this birational map is given as follows. Consider the resolution $\varpi:S\to\mathbb{P}(1,2,3)$ of $\mathbb{P}(1,2,3)$, obtained by desingularizing the $A_{1}$ and $A_{2}$ singular points, as explained in Section 2.2. The surface $S$ has a (2)(-2)-curve $C_{1}$ (i.e. a smooth rational curve of self-intersection 2-2), corresponding to the resolution of $A_{1}$, and a chain of two (2)(-2)-curves, $D_{1}$ and $D_{2}$, intersecting transversely at one point, corresponding to the resolution of $A_{2}$. The strict transform of the curve $\mathcal{\ell}$ given by $x=0$ is a rational curve $\overline{\ell}$ in $S$ of self-intersection 1-1, intersecting $C_{1}$ and $D_{1}$ at one point each, transversely. The contractions of $\overline{\ell}$ and of the transforms of $D_{1}$ and $D_{2}$, in this order, produce the projective plane, where the transform of $C_{1}$ is a straight line. See Figure 2.
@@ 3. Three special quotient foliations
We will now present the three foliations: Brunella’s very special foliation $\mathcal{F}_{3}$, and Santos’s foliations $\mathcal{F}_{4}$ and $\mathcal{F}_{6}$. The foliation $\mathcal{F}_{n}$ admits a simple description as quotient of a foliation on a rational surface under the action of a cyclic group of automorphisms of order $n$. The surface has an invariant cycle of rational curves $\Delta_{n}$ of length $n$, whose components are cyclically permuted by the automorphism, and which produces the link in the quotient. This is schematically presented in Figure 3. We will also exhibit some birational symmetries of these foliations (Theorem 5 will establish that there are no further ones).